Homework Problem: 10kg of iron (Ciron 450 Jkg C) at 200°C is put in 20kg of water (Cuater 4186 J/kg C) at 20°C. Calculate the final temperature when they reach thermal equilibrium. 7. tempfinal
Answer #1

200-T2eT-20) to)(4 18.6044( T-20) = 18-6044T-372-088 19,60YYT= S72.088